# Quantization Error Propagation: Revisiting Layer-Wise Post-Training Quantization

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive performance in various natural language processing tasks, including open-ended text generation, multi-step reasoning, and dialogue modeling. Notable examples include ChatGPT (Achiam et al., 2023) and the Llama family (Touvron et al., 2023; Grattafiori et al., 2024). However, deploying LLMs cost-effectively remains difficult because of their substantial memory usage and computational demands (Chen et al., 2023). This limitation is especially critical for edge computing and latency-sensitive applications. To address these challenges, a wide range of model compression techniques, such as quantization (Lang et al., 2024; Gong et al., 2024), pruning (Wang et al., 2024; Cheng et al., 2024), low-rank approximation (Yang et al., 2024a; Hu et al., 2022), and knowledge distillation (Xu et al., 2024a; Yang et al., 2024b), have been explored.
@@
Among these methods, layer-wise post-training quantization (PTQ) has emerged as a practical and widely used approach for large-scale LLMs (Frantar et al., 2022; Lin et al., 2024; Yao et al., 2022; Chee et al., 2023). Unlike block-wise PTQ (Tseng et al., 2024; Shao et al., 2023), global fine-tuning (Egiazarian et al., 2024; Tseng et al., 2024), quantization-aware training (QAT) (Xu et al., 2024b; Wang et al., 2023; Liu et al., 2023), and all of which require heavy retraining and backpropagation, layer-wise PTQ quantizes model parameters layer-by-layer without retraining or backpropagation, resulting in significantly lower computational and memory demands. Despite its simplicity, layer-wise PTQ effectively preserves model quality even at lower bit widths (Frantar et al., 2022; Lin et al., 2024; Chee et al., 2023). As a result, layer-wise PTQ is increasingly adopted in real-world applications due to its efficient quantization, reduced computational cost, and broader compatibility with large-scale LLMs, varying bit widths, and diverse quantization strategies.
@@
Despite significant progress in layer-wise PTQ, advancements in this area are saturating (Malinovskii et al., 2024). This study aims to push the performance boundaries of layer-wise PTQ by revisiting its core design strategy. This study begins by identifying a fundamental limitation of existing layer-wise PTQ approaches. These approaches do not adequately account for the propagation of quantization errors across layers. Quantization errors accumulate significantly, leading to a degradation in overall model performance, especially in low-bit settings. This represents a key bottleneck for the practical deployment of layer-wise PTQ in large-scale LLMs.
@@
Figure 1: WikiText-2 perplexity comparison across Llama-2 models (7B-70B) quantized to INT-4, INT-3, and INT-2, employing RTN, GPTQ, AWQ, and QuIP methods. Solid bars indicate PTQ with QEP; border bars represents PTQ without QEP. Truncated bars indicate perplexities exceeding axis limits. QEP consistently reduces perplexity, with greater improvements observed at lower bitwidths and smaller model sizes. See Section 6 for detailed settings and results.
@@
To address this issue, we propose Quantization Error Propagation (QEP), a general and computationally efficient framework that enhances the performance of layer-wise PTQ methods. QEP modifies the layer-wise optimization objective to propagate and compensate for accumulated quantization errors, while maintaining computational complexity comparable to existing layer-wise PTQ methods. Furthermore, we introduce a tunable propagation mechanism whose adjustable propagation strength prevents overfitting, a known issue previously observed in GPTQ Lin et al. (2024). This mechanism also enables adaptive control over computational overhead, especially in parameter-heavy components such as MLP blocks. Notably, the enhancement of QEP is orthogonal to existing PTQ methods and can be seamlessly integrated with any layer-wise PTQ pipeline.
@@
Extensive experiments on several LLMs across various bit-width settings show that QEP significantly enhances layer-wise PTQ methods, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024), QuIP (Chee et al., 2023), as shown in Figure 1. These improvements are particularly pronounced in extreme low-bit regimes, such as 2-bit quantization, where standard layer-wise PTQ methods typically degrade significantly.
@@ 2 Related Work
Quantization techniques primarily include data-free PTQ (Dettmers and Zettlemoyer, 2023), layer-wise PTQ (Frantar et al., 2022; Lin et al., 2024; Chee et al., 2023), block-wise PTQ (Tseng et al., 2024; Shao et al., 2023), global fine-tuning PTQ (Egiazarian et al., 2024; Tseng et al., 2024), and QAT (Xu et al., 2024b; Wang et al., 2023; Liu et al., 2023). Among these methods, weight-only layer-wise PTQ has become especially popular for large-scale models because of its computational efficiency and strong performance (Frantar et al., 2022; Lin et al., 2024; Chee et al., 2023). Recent benchmarking further highlights that most PTQ advances specifically target layer-wise methods (Zhao et al., 2025). Following the taxonomy in (Zhao et al., 2025), we outline three distinct approaches and recent developments.
@@
This category, pioneered by GPTQ (Frantar et al., 2022), uses a sequential quantization strategy, in which model weights are quantized based on the Hessian computed from a calibration dataset, while compensating for subsequent unquantized weights. Several studies refined the compensation mechanism by improving update rules (Behdin et al., 2023), integrating nonlinear quantization schemes (Liu et al., 2024a), employing adaptive grid selection (Zhang and Shrivastava, 2024), and using block-wise optimization (Guan et al., 2024).
@@
A second promising direction, advanced by QuIP (Chee et al., 2023), involves preprocessing weights through structured rotation matrices to more uniformly redistribute weight magnitudes. This approach was improved by randomized Hadamard transforms and block-wise and global fine-tuning optimization (Tseng et al., 2024). Learning-based methods to determine rotation matrices have also been introduced (Liu et al., 2024b). This rotation-based strategy has also been extended to activation quantization (Ashkboos et al., 2024).
@@
Other approaches focus on identifying and preserving salient weights, often using mixed-precision quantization frameworks (Dettmers et al., 2022, 2023; Shang et al., 2023). Although mixed-precision methods usually add complexity due to various data types, AWQ (Lin et al., 2024) mitigates these implementation difficulties. AWQ strategically employs a global scaling mechanism to align salient weights with the quantization grid better, simplifying deployment while maintaining high accuracy.
Recent advances in layer-wise PTQ have mainly focused on nonlinear quantization and block-wise and global fine-tuning extensions; however, the fundamental layer-wise optimization has remained largely unchanged since GPTQ (Frantar et al., 2022). This study revisits this foundational strategy, identifies its key limitations, and proposes improvements, demonstrating performance gains on the fundamental benchmarks such as GPTQ (Frantar et al., 2022), QuIP (Chee et al., 2023), and AWQ (Lin et al., 2024). Therefore, our contributions complement and are orthogonal to recent advancements, such as nonlinear quantization and structured extensions.
@@ 3 Background
Layer-wise PTQ has emerged as a promising framework (Frantar et al., 2022; Frantar and Alistarh, 2022) for compressing large-scale LLMs. Recent advancements in this area have significantly reduced the computational overhead and memory requirements of deploying LLMs. Despite methodological differences, existing layer-wise PTQ approaches typically follow a shared sequential quantization scheme, processing each layer independently and sequentially from the input layer toward the output layer.
@@
Existing PTQ methods typically use one of two possible forms for activation inputs $\mathsf{X}_{l}$: Either quantized activations ${\bm{X}}_{l}$, obtained by forward propagating the calibration dataset through previously quantized weights $\{\widehat{{\bm{W}}}_{1},\ldots,\widehat{{\bm{W}}}_{l-1}\}$, or full-precision activations ${\bm{X}}_{l}$, resulting from forward propagation through the original, unquantized weights $\{{\bm{W}}_{1},\ldots,{\bm{W}}_{l-1}\}$. There is no consensus among existing PTQ methods (Frantar et al., 2022; Lin et al., 2024; Chee et al., 2023) regarding whether quantized or full-precision activations produce better quantization outcomes.
Leading layer-wise PTQ methods use distinct optimization strategies to approximate the behavior of the original model while adhering to the foundational sequential layer-wise framework in Eq. (1). GPTQ (Frantar et al., 2022), for example, uses quantized activations, $\mathsf{X}_{l}=\widehat{{\bm{X}}}_{l}$, and quantizes parameters row-wise by sequentially minimizing reconstruction error and correcting residuals in the remaining unquantized entries until each row is fully quantized. AWQ Lin et al. (2024) uses original activations, $\mathsf{X}_{l}={\bm{X}}_{l}$, and identifies a small subset of salient weights whose magnitudes significantly influence the layer outputs, subsequently rescaling these weights before quantization.
@@ 4 Bottleneck: Quantization Error Accumulation and Growth
Figure 2: Accumulation and growth of quantization errors across layers in a partially quantized Llama2-7B model (Touvron et al., 2023). The first $10$ Transformer blocks are quantized using standard RTN (BASE) and QEP-enhanced RTN (With QEP), while the remaining Transformer blocks after the $10$th remain at full precision. The plot shows the squared Frobenius norm $\Delta_{m}$, defined in Eq. (2), between the original and partially quantized outputs at each Transformer block $m$.
To motivate our proposed approach, this section first revisits the core layer-wise optimization formulation given by Eq. (1), emphasizing its key limitation: The accumulation and growth of quantization errors across layers significantly degrade the performance.
We investigate this phenomenon using experiments conducted on the pre-trained Llama-2-7B model (Touvron et al., 2023). Specifically, we quantize only the first $10$ Transformer blocks (Vaswani et al., 2017), while keeping all subsequent blocks in full precision. To quantify the propagation and accumulation of the errors, we measure the discrepancy between fully precise and partially quantized outputs at each block using a calibration dataset. Let $\mathtt{TransBlock}_{m}(\cdot)$ denote the original full-precision $m$-th Transformer block, and $\widehat{\mathtt{TransBlock}}_{m}(\cdot)$ denote its quantized counterpart. We evaluate the following metric at the $m$-th block:
@@
where
$f_{m}({\bm{X}})\coloneqq\mathtt{TransBlock}_{m}\circ\cdots\circ\mathtt{TransBlock}_{n+1}\circ\mathtt{TransBlock}_{n}\circ\cdots\circ\mathtt{TransBlock}_{1}({\bm{X}}),$
$\widehat{f}_{m}({\bm{X}})\coloneqq\mathtt{TransBlock}_{m}\circ\cdots\circ\mathtt{TransBlock}_{n+1}\circ\widehat{\mathtt{TransBlock}}_{n}\circ\cdots\circ\widehat{\mathtt{TransBlock}}_{1}({\bm{X}}).$
This experiment sets $n=10$. Figure 2 shows an approximately exponential accumulation or errors within the quantized layer, as well as an error growth that persists in the unquantized layers. This growth occurs due to the layer-wise independent quantization approach described in Eq. (1), which neither accounts for quantization error propagated from previous layers nor corrects previously accumulated errors, thus exacerbating error growth in subsequent unquantized layers. The exponential accumulation of quantization errors observed empirically can also be theoretically explained under mild conditions, as detailed in Appendix A.2. Therefore, instead of treating layer-wise quantization as a series of independent optimization problems, it is essential to reformulate the original layer-wise optimization presented in Eq. (1) to mitigate error accumulation and growth.
@@ 5 QEP: Quantization Error Propagation
Existing layer-wise independent PTQ has inherent limitations, particularly the accumulation and growth of quantization errors discussed in 4. To address these limitations, we introduce Quantization Error Propagation (QEP), a general, lightweight, and scalable framework that improves layer-wise PTQ by propagating quantization errors. In subsequent sections, we provide theoretical evidence showing that QEP effectively reduces quantization errors.
@@ 5.2 Weight Correction
The proof of Proposition 5.1 is provided in Appendix A.1. Proposition 5.1 highlights an important distinction from the existing formulation given by Eq. (1). Specifically, when upstream quantization introduces non-negligible errors, i.e., ${\bm{\delta}}_{l-1}\neq 0$, the optimal quantized weights differ from straightforward approximations of the original weights ${\bm{W}}_{l}$. Instead, the optimal solution explicitly includes a correction term that compensating for accumulated quantization errors.
@@
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\left\|{\bm{W}}_{l}^{\ast}\widehat{{\bm{X}}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\right\|_{F}^{2},$ (5)
@@
$f_{{\bm{\theta}}}(X)=\sigma_{L}({\bm{W}}_{L}\sigma_{L-1}({\bm{W}}_{L-1}\cdots\sigma_{2}({\bm{W}}_{2}\sigma_{1}({\bm{W}}_{1}{\bm{X}}))\cdots)).$ (6)
@@
Explicit conditions and detailed proof are provided in Appendix A.3. The additional computational overhead arises solely from computing the correction term ${\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$, since computing the Hessian inverse $\widehat{{\bm{H}}}_{l}^{-1}$ remains unchanged from existing layer-wise independent PTQ. As empirically demonstrated in Section 6.1, this additional computation requires significantly less runtime compared to the quantization processes of layer-wise PTQ methods, even for large-scale LLMs, due to the tunable mechanism described in the next section.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive performance in various natural language processing tasks, including open-ended text generation, multi-step reasoning, and dialogue modeling. Notable examples include ChatGPT (achiam2023gpt) and the Llama family (touvron2023llama; grattafiori2024llama). However, deploying LLMs cost-effectively remains difficult because of their substantial memory usage and computational demands (chen2023frugalgpt). This limitation is especially critical for edge computing and latency-sensitive applications. To address these challenges, a wide range of model compression techniques, such as quantization (lang2024comprehensive; gong2024survey), pruning (wang2024model; cheng2024survey), low-rank approximation (yang2024low; hu2022lora), and knowledge distillation (xu2024survey; yang2024survey), have been explored.
@@
Among these methods, layer-wise post-training quantization (PTQ) has emerged as a practical and widely used approach for large-scale LLMs (frantar2022gptq; lin2024awq; yao2022zeroquant; chee2023quip). Unlike block-wise PTQ (tseng2024quip; shao2023omniquant), global fine-tuning (egiazarian2024extreme; tseng2024quip), quantization-aware training (QAT) (xu2024onebit; wang2023bitnet; liu2023llm), and all of which require heavy retraining and backpropagation, layer-wise PTQ quantizes model parameters layer-by-layer without retraining or backpropagation, resulting in significantly lower computational and memory demands. Despite its simplicity, layer-wise PTQ effectively preserves model quality even at lower bit widths (frantar2022gptq; lin2024awq; chee2023quip). As a result, layer-wise PTQ is increasingly adopted in real-world applications due to its efficient quantization, reduced computational cost, and broader compatibility with large-scale LLMs, varying bit widths, and diverse quantization strategies.
@@
Despite significant progress in layer-wise PTQ, advancements in this area are saturating (malinovskii2024pv). This study aims to push the performance boundaries of layer-wise PTQ by revisiting its core design strategy. This study begins by identifying a fundamental limitation of existing layer-wise PTQ approaches. These approaches do not adequately account for the propagation of quantization errors across layers. Quantization errors accumulate significantly, leading to a degradation in overall model performance, especially in low-bit settings. This represents a key bottleneck for the practical deployment of layer-wise PTQ in large-scale LLMs.
@@
Figure 1: WikiText‑2 perplexity comparison across Llama‑2 models (7B-70B) quantized to INT‑4, INT‑3, and INT‑2, employing RTN, GPTQ, AWQ, and QuIP methods. Solid bars indicate PTQ with QEP; border bars represent PTQ without QEP. Truncated bars indicate perplexities exceeding axis limits. QEP consistently reduces perplexity, with greater improvements observed at lower bitwidths and smaller model sizes. See Section 6 for detailed settings and results.
@@
To address this issue, we propose Quantization Error Propagation (QEP), a general and computationally efficient framework that enhances the performance of layer-wise PTQ methods. QEP modifies the layer-wise optimization objective to propagate and compensate for accumulated quantization errors, while maintaining computational complexity comparable to existing layer-wise PTQ methods. Furthermore, we introduce a tunable propagation mechanism whose adjustable propagation strength prevents overfitting, a known issue previously observed in GPTQ (lin2024awq). This mechanism also enables adaptive control over computational overhead, especially in parameter-heavy components such as MLP blocks. Notably, the enhancement of QEP is orthogonal to existing PTQ methods and can be seamlessly integrated with any layer-wise PTQ pipeline.
@@
Extensive experiments on several LLMs across various bit-width settings show that QEP significantly enhances layer-wise PTQ methods, including GPTQ (frantar2022gptq), AWQ (lin2024awq), QuIP (chee2023quip), as shown in Figure 1. These improvements are particularly pronounced in extreme low-bit regimes, such as 2-bit quantization, where standard layer-wise PTQ methods typically degrade significantly.
@@ 2 Related Work
Quantization techniques primarily include data-free PTQ (dettmers2023case), layer-wise PTQ (frantar2022gptq; lin2024awq; chee2023quip), block-wise PTQ (tseng2024quip; shao2023omniquant), global fine-tuning PTQ (egiazarian2024extreme; tseng2024quip), and QAT (xu2024onebit; wang2023bitnet; liu2023llm). Among these methods, weight-only layer-wise PTQ has become especially popular for large-scale models because of its computational efficiency and strong performance (frantar2022gptq; lin2024awq; chee2023quip). Recent benchmarking further highlights that most PTQ advances specifically target layer-wise methods (zhao2025benchmarking). Following the taxonomy in (zhao2025benchmarking), we outline three distinct approaches and recent developments.
@@
This category, pioneered by GPTQ (frantar2022gptq), uses a sequential quantization strategy, in which model weights are quantized based on the Hessian computed from a calibration dataset, while compensating for subsequent unquantized weights. Several studies refined the compensation mechanism by improving update rules (behdin2023quantease), integrating nonlinear quantization schemes (liu2024vptq), employing adaptive grid selection (zhang2024leanquant), and using block-wise optimization (guan2024aptq).
@@
A second promising direction, advanced by QuIP (chee2023quip), involves preprocessing weights through structured rotation matrices to more uniformly redistribute weight magnitudes. This approach was improved by randomized Hadamard transforms and block-wise and global fine-tuning optimization (tseng2024quip). Learning-based methods to determine rotation matrices have also been introduced (liu2024spinquant). This rotation-based strategy has also been extended to activation quantization (ashkboos2024quarot).
@@
Other approaches focus on identifying and preserving salient weights, often using mixed-precision quantization frameworks (dettmers2022gpt3; dettmers2023spqr; shang2023pb). Although mixed-precision methods usually add complexity due to various data types, AWQ (lin2024awq) mitigates these implementation difficulties. AWQ strategically employs a global scaling mechanism to align salient weights with the quantization grid better, simplifying deployment while maintaining high accuracy.
Recent advances in layer-wise PTQ have mainly focused on nonlinear quantization and block-wise and global fine-tuning extensions; however, the fundamental layer-wise optimization has remained largely unchanged since GPTQ (frantar2022gptq). This study revisits this foundational strategy, identifies its key limitations, and proposes improvements, demonstrating performance gains on the fundamental benchmarks such as GPTQ (frantar2022gptq), QuIP (chee2023quip), and AWQ (lin2024awq). Therefore, our contributions complement and are orthogonal to recent advancements, such as nonlinear quantization and structured extensions.
@@ 3 Background
Layer-wise PTQ has emerged as a promising framework (frantar2022gptq; frantar2022optimal) for compressing large-scale LLMs. Recent advancements in this area have significantly reduced the computational overhead and memory requirements of deploying LLMs. Despite methodological differences, existing layer-wise PTQ approaches typically follow a shared sequential quantization scheme, processing each layer independently and sequentially from the input layer toward the output layer.
@@
Existing PTQ methods typically use one of two possible forms for activation inputs $\mathsf{X}_{l}$: Either quantized activations ${\bm{X}}_{l}$, obtained by forward propagating the calibration dataset through previously quantized weights $\{\widehat{{\bm{W}}}_{1},\ldots,\widehat{{\bm{W}}}_{l-1}\}$, or full-precision activations ${\bm{X}}_{l}$, resulting from forward propagation through the original, unquantized weights $\{{\bm{W}}_{1},\ldots,{\bm{W}}_{l-1}\}$. There is no consensus among existing PTQ methods (frantar2022gptq; lin2024awq; chee2023quip) regarding whether quantized or full-precision activations produce better quantization outcomes.
Leading layer-wise PTQ methods use distinct optimization strategies to approximate the behavior of the original model while adhering to the foundational sequential layer-wise framework in Eq. (1). GPTQ (frantar2022gptq), for example, uses quantized activations, $\mathsf{X}_{l}=\widehat{{\bm{X}}}_{l}$, and quantizes parameters row-wise by sequentially minimizing reconstruction error and correcting residuals in the remaining unquantized entries until each row is fully quantized. AWQ (lin2024awq) uses original activations, $\mathsf{X}_{l}={\bm{X}}_{l}$, and identifies a small subset of salient weights whose magnitudes significantly influence the layer outputs, subsequently rescaling these weights before quantization.
@@ 4 Bottleneck: Quantization Error Accumulation and Growth
Figure 2: Accumulation and growth of quantization errors across layers in a partially quantized Llama2-7B model (touvron2023llama). The first $10$ Transformer blocks are quantized using standard RTN (BASE) and QEP-enhanced RTN (With QEP), while the remaining Transformer blocks after the 1010th remain at full precision. The plot shows the squared Frobenius norm $\Delta_{m}$, defined in Eq. (2), between the original and partially quantized outputs at each Transformer block $m$.
To motivate our proposed approach, this section first revisits the core layer-wise optimization formulation given by Eq. (1), emphasizing its key limitation: The accumulation and growth of quantization errors across layers significantly degrade the performance. We investigate this phenomenon using experiments conducted on the pre-trained Llama-2-7B model (touvron2023llama). Specifically, we quantize only the first $10$ Transformer blocks (vaswani2017attention), while keeping all subsequent blocks in full precision. To quantify the propagation and accumulation of the errors, we measure the discrepancy between fully precise and partially quantized outputs at each block using a calibration dataset. Let $\mathtt{TransBlock}_{m}(\cdot)$ denote the original full-precision $m$-th Transformer block, and $\widehat{\mathtt{TransBlock}}_{m}(\cdot)$ denote its quantized counterpart. We evaluate the following metric at the $m$-th block:
@@
$f_{m}({\bm{X}})\coloneqq\mathtt{TransBlock}_{m}\circ\cdots\circ\mathtt{TransBlock}_{n+1}\circ\mathtt{TransBlock}_{n}\circ\cdots\circ\mathtt{TransBlock}_{1}({\bm{X}}),$ (3)
$\widehat{f}_{m}({\bm{X}})\coloneqq\mathtt{TransBlock}_{m}\circ\cdots\circ\mathtt{TransBlock}_{n+1}\circ\widehat{\mathtt{TransBlock}}_{n}\circ\cdots\circ\widehat{\mathtt{TransBlock}}_{1}({\bm{X}}).$ (4)
This experiment sets $n=10$. Figure 2 shows an approximately exponential accumulation or errors within the quantized layer, as well as an error growth that persists in the unquantized layers. This growth occurs due to the layer-wise independent quantization approach described in Eq. (1), which neither accounts for quantization error propagated from previous layers nor corrects previously accumulated errors, thus exacerbating error growth in subsequent unquantized layers. The exponential accumulation of quantization errors observed empirically can also be theoretically explained under mild conditions, as detailed in Appendix B.2. Therefore, instead of treating layer-wise quantization as a series of independent optimization problems, it is essential to reformulate the original layer-wise optimization presented in Eq. (1) to mitigate error accumulation and growth.
@@ 5 QEP: Quantization Error Propagation
Existing layer-wise independent PTQ has inherent limitations, particularly the accumulation and growth of quantization errors discussed in Section 4. To address these limitations, we introduce Quantization Error Propagation (QEP), a general, lightweight, and scalable framework that improves layer-wise PTQ by propagating quantization errors. In subsequent sections, we provide theoretical evidence showing that QEP effectively reduces quantization errors.
@@ 5.2 Weight Correction
The proof of Proposition 5.1 is provided in Appendix B.1. Proposition 5.1 highlights an important distinction from the existing formulation given by Eq. (1). Specifically, when upstream quantization introduces non-negligible errors, i.e., ${\bm{\delta}}_{l-1}\neq 0$, the optimal quantized weights differ from straightforward approximations of the original weights ${\bm{W}}_{l}$. Instead, the optimal solution explicitly includes a correction term that compensating for accumulated quantization errors.
@@
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\left\|{\bm{W}}_{l}^{\ast}\widehat{{\bm{X}}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\right\|_{F}^{2}.$ (7)
@@
$f_{{\bm{\theta}}}(X)=\sigma_{L}({\bm{W}}_{L}\sigma_{L-1}({\bm{W}}_{L-1}\cdots\sigma_{2}({\bm{W}}_{2}\sigma_{1}({\bm{W}}_{1}{\bm{X}}))\cdots)),$ (8)
@@
Explicit conditions and detailed proof are provided in Appendix B.3. The additional computational overhead arises solely from computing the correction term ${\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$, since computing the Hessian inverse $\widehat{{\bm{H}}}_{l}^{-1}$ remains unchanged from existing layer-wise independent PTQ. As empirically demonstrated in Section 6.1, this additional computation requires significantly less runtime compared to the quantization processes of layer-wise PTQ methods, even for large-scale LLMs, due to the tunable mechanism described in the next section.